Two balls are drawn in succession without replacement of an containing 5 white balls and 6 black balls Let B be the random variable representing the number of black balls. Construct the probability distribution of the random variable R


1
Expert's answer
2022-01-31T16:03:49-0500

Let "X=" the random variable representing the number of black balls in result.

There are "5+6=11" balls. Then


"P(W)=5\/11,P(B)=6\/11"


"S=\{WW, WB, BW, BB\}"

The posiible value of "X: 0, 1, 2."


"P(X=0)=P(WW)=\dfrac{\dbinom{6}{0}\dbinom{5}{2-0}}{\dbinom{11}{2}}"

"=\dfrac{1(10)}{55}=\dfrac{2}{11}"


"P(X=1)=P(WB)+P(BW)=\dfrac{\dbinom{6}{1}\dbinom{5}{2-1}}{\dbinom{11}{2}}"

"=\dfrac{6(5)}{55}=\dfrac{6}{11}"


"P(X=2)=P(BB)=\dfrac{\dbinom{6}{2}\dbinom{5}{2-2}}{\dbinom{11}{2}}"

"=\dfrac{15(1)}{55}=\dfrac{3}{11}"



Construct the probability distribution of the random variable "X"


"\begin{matrix}
 x & 0 & 1 & 2 \\
 p(x) & 2/11 & 6/11 & 3/11
\end{matrix}"